# A Probabilistic Framework for LLM Hallucination Detection via Belief Tree Propagation

### v1

@@ 3.4 A Hidden Markov Tree Model
Figure 3 shows an example hidden Markov tree model built upon the belief tree, where there are two layers of variables. The upper layer consists of the confidence score of each statement, i.e., $\{S_{u}\}$, which is estimated during the belief tree construction process. The lower layer consists of the binary variables representing the actual correctness of each statement, denoted as $\{Z_{u}\}$. $Z_{u}=1$ if statement $v$ is correct and $Z_{u}=0$ otherwise. $\{S_{u}\}$ are observed variables while $\{Z_{u}\}$ are hidden variables.
Given the above hidden Markov tree model, determining the truthfulness of the target statement can be cast as computing the posterior probability, i.e., $p(Z_{0}=1|\{S_{u}\})$, which means the truthfulness of the root node given the confidence scores on all the nodes in this belief tree. According to the probability graph in figure 3, we need to estimate the following probabilities.
@@
The first set is the emission probability, $p(S_{u}|Z_{u})$, which characterizes the LLM’s confidence score of the statement given its underlying truthfulness. The emission probability captures the quality of LLM’s original confidence score. If the confidence score is accurate, then $p(S_{u}|Z_{u}=1)$ will concentrate its probability mass towards $1$, and $p(S_{u}|Z_{u}=0)$ will concentrate its probability mass towards 00. Otherwise, the conditional distributions will be more spread.
The second set of distribution is the transition probability, $p(Z_{\mathcal{C}(u)}|Z_{u})$, where $Z_{\mathcal{C}(u)}$ represent the set of truthfulness variables of all the child nodes of $u$. The transition probability is the joint probability of the truthfulness of the child statements given that of the parent statement, which captures the logical relationships among these statements.
The third set of distribution is the prior distribution of hidden variables, i.e., $p(Z_{u}=z)$, which refers to the probability of an arbitrary statement $u$ generated by the LLM being true or false. We use uniform prior (both 0.5 for $z=1$ and $z=0$) in this paper.
Therefore, the process of computing the posterior probability contains two steps: ❷ the estimation step, to first estimate these probabilities on a held-out dataset. ❸ the inference step, to compute the posterior probability for each testing example. Section 3.5 will discuss how to determine the emission and transition probabilities and Section 3.6 how to infer $p(Z_{0}=1|\{S_{u}\})$.
@@ 3.5 Determining Conditional Probabilities
The emission probabilities, $p(S_{u}|Z_{u})$, can be estimated from labeled datasets. Specifically, given a dataset of statements with truthfulness labels, we can run the LLM to obtain the confidence scores of all the statements. We can then estimate $p(S_{u}|Z_{u}=1)$ by obtaining the empirical distribution of the confidence scores within the statements labeled as true, and $p(S_{u}|Z_{u}=0)$ within the statements labeled as false. To obtain the empirical distribution of the continuous confidence scores, we first quantize the support $[0,1]$ into multiple bins, and count the histogram of the confidence scores falling into each bin. The boundaries of each bin are listed in Table 1.
Table 1: Empirical estimation of the emission probability on Wikibio-GPT3 dataset [4].
@@
The transition probability, $p(Z_{\mathcal{C}(u)}|Z_{u})$, is a multivariate Bernoulli distribution with the support $\{0,1\}^{m}$, where $m$ is the size of $\mathcal{C}(u)$. As discussed in Section 3.3, different statement generation strategies have different logical relationship properties. For the statement decomposition strategy, the child statements are always jointly equivalent to the parent statement. Therefore, when the parent claim is true, i.e., $Z_{u}=1$, we know that all of the child statements must be true, and thus $p(Z_{\mathcal{C}(u)}=\bm{z}|Z_{u}=1)$ equals one when $\bm{z}$ is an all-one vector and zero when $\bm{z}$ takes other values. Conversely, when the parent claim is false, i.e., $Z_{u}=0$, we can infer that at least one child statement is wrong, so $p(Z_{\mathcal{C}(u)}=\bm{z}|Z_{u}=0)$ is set to zero when $\bm{z}$ is an all-one vector and uniformly set to $1/(2^{m}-1)$ when $\bm{z}$ takes other values.
For the other two strategies, the child statements of the same parent statement are considered independent conditional on the parent statement. Therefore, we can decompose the probability $p(Z_{\mathcal{C}(u)}|Z_{u})$ into $\prod_{v\in\mathcal{C}(u)}p(Z_{v}|Z_{u})$, where each $p(Z_{v}|Z_{u})$ is the transitional probability between parent statement $u$ and one child statement $v$. $p(Z_{v}|Z_{u})$ is determined based on the four different possible logical relationships between $u$ and $v$ (recall the discussion in Section 3.3), which is listed in Table 2, where $p_{t}$ and $p_{f}$ are two hyper-parameters.
@@
$u\Leftrightarrow v$ $u\Rightarrow v$ $u\Leftarrow v$ $u\Rightarrow\neg v$
$u=\texttt{True}$ $u=\texttt{False}$ $u=\texttt{True}$ $u=\texttt{False}$ $u=\texttt{True}$ $u=\texttt{False}$ $u=\texttt{True}$ $u=\texttt{False}$
$v=\texttt{True}$ 1.0 0.0 1.0 $p_{t}$ $p_{t}$ 0.0 0.0 $p_{t}$
$v=\texttt{False}$ 0.0 1.0 0.0 $p_{f}$ $p_{f}$ 1.0 1.0 $p_{f}$
@@ 3.6 Inferring the Posterior Probability of $Z_{0}$
Inferring the posterior probability $p(Z_{0}=1|\{S_{u}\})$ is a standard inference problem in hidden Markov tree models and we can directly apply the standard upward-downward algorithm [16] to efficiently compute this probability. For brevity, we will only describe the gist of the algorithm here. More details can be found in Appendix A.3 and previous work [16, 17].
The algorithm introduces an auxiliary conditional distribution $\beta(z,u)\equiv p(S_{\mathcal{T}(u)}|Z_{u}=z)$, where $z\in\{0,1\}$ and $S_{\mathcal{T}(u)}$ represents the confidence scores of all the nodes in the sub-tree whose root node is $u$. By definition, the posteior probability of our interest, $p(Z_{0}=1|\{S_{\mathcal{T}(0)}\})$, can be represented as $\beta(z=1,0)/(\beta(z=0,0)+\beta(z=1,0))$. According to the Bayesian rule, we can derive a recursive relationship between $\beta(z,u)$ of a parent node $u$ and those of the child nodes:
$\small\beta(z,u)=p(S_{u}|Z_{u}=z)\cdot\sum_{Z_{\mathcal{C}(u)}\in\{0,1\}^{m}}p(Z_{\mathcal{C}(u)}|Z_{u}=z)\prod_{v\in\mathcal{C}(u)}\beta(Z_{v},v),$ (1)
where $m$ is the size of $\mathcal{C}(u)$ (namely the number of child statements to node $u$). Note that the first term on the RHS is the emission probability, and the first term inside the summation is the transition probability, which are both already known. Therefore, Equation 1 provides a way to compute $\beta(z,0)$ recursively from the leaf nodes back to the root node. First, we compute the $\beta(z,u)$ of the leaf nodes as $\beta(z,u)=p(S_{u}|Z_{u}=z)$. Then, we use Equation 1 to recursively compute the $\beta(z,u)$ of a parent node $u$ from their child nodes, until the root node is reached. The entire process is essentially propagating and merging the beliefs in the confidence scores from sub-trees upward to the parent node. By the time we reach the root node, we have gathered the information of all the confidence scores. Hence this process is also referred to as a belief propagation process.
@@ 4 Experiments
In this section, we conduct empirical evaluations on commonly-used hallucination detection benchmarks to demonstrate the effectiveness of the proposed method.
@@ 4.1 Experiment Configurations
We follow the previous work of hallucination detection [4, 15] and use the following datasets for evaluation: Wikibio-GPT3 [4], FELM-Science [15], and FactCheckGPT [43]. Wikibio-GPT3 mainly consists of biography articles generated by LLMs, whereas FELM-Science and FactCheckGPT cover a broader range of topics such as physics, chemistry, and computer science.
Evaluation settings and metrics
Following the default settings on each dataset, we conduct hallucination detection at sentence-level on Wikibio-GPT3 and FactCheckGPT and segment-level on FELM-Science. Our evaluation encompasses a comprehensive set of metrics, including the area under the receiver operator characteristic curve (AUROC), area under the precision-recall curve (AUC-PR), F1 score, and detection accuracy. As F1 score and detection accuracy evaluation require a decision threshold, we search for the optimal threshold to maximize the F1 score for each method and compute the two metrics. Note that when computing these metrics, we consider hallucinated examples as positive instances, following the default configurations of these datasets. Moreover, Wikibio-GPT3 and FELM-Science exhibit significant class imbalances, so the detection accuracy on these datasets serve merely as reference points.
Baselines
We include the following baselines for comparison. The implementation details are available in Appendix A.1. (1) Prior confidence, which directly queries the model’s confidence regarding the truthfulness of each sentence or segment. (2) Chain-of-thought, which prompts the model to first generate a reasoning process before deciding the truthfulness. We adopt the Chain-of-thought prompting method from the official FELM [15] dataset, with slight modifications for sentence-level and segment-level hallucination detection. (3) SelfCheckGPT [4], which samples additional responses from the model and use the inconsistency between each response and the target statement for hallucination detection. Among the multiple variants, we choose SelfCheckGPT-prompt for comparison, which achieves the best performance. We sample 20 responses for hallucination detection following its default configuration.(4) Maieutic-Prompting [8], which first builds a belief tree via backward-chaining and then infers the truth-value of the original statement that resolves the inconsistencies.
Implementation details
We evaluate our method along with baselines using both closed-sourced and open-sourced models, including GPT-3.5-turbo-0125 and Llama-3-8B-Instruct. For our method, we set the maximum belief tree depth to $2$. We employ greedy decoding during belief tree construction and prior belief estimation. The exception is the statement correction strategy, where we sample $5$ corrected statements using temperature $0.7$. Additionally, since each statement in FactCheckGPT is manually processed to ensure it contains only one property or fact, we do not apply statement decomposition when build the belief tree for statements from FactCheckGPT. We use the first 120 examples in the Wikibio-GPT3 dataset to estimate the emission probability in our method, and validate it on the remaining examples and two other datasets. More implementation details can be found in Appendix A.1.
4.2 Experiment Results
Table 3: Hallucination detection performance of different methods. We report AUROC, ROC-PR, F1 score, and detection accuracy(Acc) for all methods with two different backbone models. The best results are highlighted in bold.
@@ 4.2 Experiment Results
We evaluate the effectiveness of BTProp with the experiment results in Table 3. We highlight the following observations. First, our method achieves the best performance on FELM-Science and FactCheckGPT datasets across different backbones, demonstrating the superiority of our method. BTProp improves upon the best baselines by 3% - 9% on AUROC and ROC-PR. The only exception is the Wikibio-GPT3 dataset, where the SelfCheckGPT is more effective for detection hallucinated outputs in biographies generated by LLMs. Second, compared to SelfCheckGPT which leverages contradictions between the target statement and the sampled responses for hallucination detection, our method is more effective on detecting hallucinated responses related to scientific knowledge. Both FELM-Science and FactCheckGPT datasets contain a significant proportion of questions on scientific knowledge, and our method achieves the best performance on them. Third, chain-of-thought prompting is less effective in hallucination detection, especially on the FELM-Science dataset. This finding aligns with the experimental results in the original FELM dataset paper. Our experiments show that the model tends to regard the input sentence as true most of the time, leading to sub-optimal performance. Finally, the Maieutic-prompting method, which is originall designed to verify the correctness of statements related to commonsense reasoning, is less effective for hallucination detection. Even on the FELM-Science and FactCheckGPT datasets which contain many statements about scientific knowledge, it remains less effective compared to other methods.
@@
One concern on our method is the belief tree construction would be time-consuming and inefficient. Therefore, we compare the time cost of our method with the baselines on the FELM-Science and FactCheckGPT datasets, where our method significantly outperform baselines. We use the open-sourced Llama-3-8b-Instruct model as the backbone, and eliminate CoT prompting method on FELM-Science dataset due to its poor performance. We visualize the trade-off between hallucination detection performance and the time cost in Figure 4 of different methods. The baseline methods, including SelfCheckGPT and CoT prompting, samples additional 20 responses for hallucination detection by default. We vary the number of samples and test the corresponding performance to visualize the trade-off. As depicted in the graph, the baseline performance shows diminishing returns in performance improvement with increased computational complexity. As time cost increases, the performance growth of the two baseline models gradually levels off, eventually reaching a point where additional time does not translate into significantly better performance. In comparison, our method can effectively trades off increased time cost for significantly improved performance. Despite its higher computational complexity, it achieves a marked improvement in performance metrics, underscoring its efficiency and effectiveness in hallucination detection compared to the baseline models.
@@
We visualize several examples of the constructed belief trees to demonstrate how the inconsistencies in model’s beliefs help detect the hallucination. Due to space limit, we put these examples and the corresponding analyses in Appendix A.2.

### v2

@@ 3.3 A Hidden Markov Tree Model
Figure 3 shows an example hidden Markov tree model built upon the belief tree, where there are two layers of variables. The upper layer consists of the confidence score of each statement, denoted $\{S_{u}\}$, which is estimated during the belief tree construction process. The lower layer consists of the binary variables representing the actual correctness of each statement, denoted as $\{Z_{u}\}$. $Z_{u}=T$ if statement $v$ is correct and $Z_{u}=F$ otherwise. $\{S_{u}\}$ are observed variables while $\{Z_{u}\}$ are hidden variables.
Given the above hidden Markov tree model, determining the truthfulness of the target statement can be cast as computing the posterior probability, i.e., $p(Z_{0}=T\mid\{S_{u}\})$, which means the truthfulness of the root node given the confidence scores on all the nodes in this belief tree. To computing $p(Z_{0}=T\mid\{S_{u}\})$, we need to estimate the following probabilities.
@@
The first set is the emission probability, $p(S_{u}\mid Z_{u})$, which characterizes the LLM’s confidence score of the statement given its underlying truthfulness. The emission probability captures the quality of LLM’s original confidence score. If the confidence score is accurate, then $p(S_{u}\mid Z_{u}=T)$ will concentrate its probability mass towards $1$, and $p(S_{u}\mid Z_{u}=F)$ will concentrate its probability mass towards 00. Otherwise, the conditional distributions will be more spread.
The second set of distribution is the transition probability, $p(Z_{\mathcal{C}(u)}\mid Z_{u})$, where $Z_{\mathcal{C}(u)}$ represent the set of truthfulness variables of all the child nodes of $u$. The transition probability is the joint probability of the truthfulness of the child statements given that of the parent statement, which captures the logical relationships among these statements.
The third set of distribution is the prior distribution of hidden variables, i.e., $p(Z_{u}=z)$, which refers to the probability of an arbitrary statement $u$ generated by the LLM being true or false. We use uniform prior (both 0.5 for $z=T$ and $z=F$) in this paper.
Therefore, the process of computing the posterior probability contains two steps: ❷ the estimation step, to first estimate these probabilities on a held-out dataset. ❸ the inference step, to compute the posterior probability for each testing example. Section 3.4 will discuss how to determine the emission and transition probabilities and Section 3.5 how to infer $p(Z_{0}=T\mid\{S_{u}\})$.
@@ 3.4 Determining Conditional Probabilities
The emission probabilities, $p(S_{u}\mid Z_{u})$, can be estimated from labeled datasets. Specifically, given a dataset of statements with truthfulness labels, we can run the LLM to obtain the confidence scores of all the statements. We can then estimate $p(S_{u}\mid Z_{u}=T)$ by obtaining the empirical distribution of the confidence scores within the statements labeled as true, and $p(S_{u}\mid Z_{u}=F)$ within the statements labeled as false. To obtain the empirical distribution of the continuous confidence scores, we first quantize the support $[0,1]$ into multiple bins, and count the histogram of the confidence scores falling into each bin. The boundaries of each bin are listed in Table 1. Given a particular confidence score on a statement, we look up the table to get the corresponding emission probability. For example, a confidence score 0.95 leads to $p(S_{u}=0.95\mid Z_{u}=F)=0.65$.
@@
Table 1: Empirical estimation of the emission probability on Wikibio-GPT3 dataset Manakul et al. (2023).
@@
The transition probability, $p(Z_{\mathcal{C}(u)}\mid Z_{u})$, is a multivariate Bernoulli distribution with the support $\{T,F\}^{m}$, where $m$ is the size of $\mathcal{C}(u)$. As discussed in Section 3.2, different statement generation strategies have different logical relationship properties. For the statement decomposition strategy, the child statements are always jointly equivalent to the parent statement. Therefore, when the parent claim is true, i.e., $Z_{u}=T$, we know that all of the child statements must be true, and thus $p(Z_{\mathcal{C}(u)}=\bm{z}\mid Z_{u}=T)$ equals one when every element in $\bm{z}$ is true and zero otherwise. Conversely, when the parent claim is false, i.e., $Z_{u}=F$, we can infer that at least one child statement is wrong, so $p(Z_{\mathcal{C}(u)}=\bm{z}\mid Z_{u}=F)$ is set to zero when every element in $\bm{z}$ is true and uniformly set to $1/(2^{m}-1)$ otherwise.
For the other two strategies, the child statements of the same parent statement are considered independent conditional on the parent statement. Therefore, we can decompose the probability $p(Z_{\mathcal{C}(u)}\mid Z_{u})$ into $\prod_{v\in\mathcal{C}(u)}p(Z_{v}\mid Z_{u})$, where each $p(Z_{v}\mid Z_{u})$ is the transitional probability between parent statement $u$ and one child statement $v$. $p(Z_{v}\mid Z_{u})$ is determined based on the four different possible logical relationships between $u$ and $v$ (recall the discussion in Section 3.2), which is listed in Table 2.
$u\Leftrightarrow v$ $u\Rightarrow v$ $u\Leftarrow v$ $u\Rightarrow\neg v$
$u=T$ $u=F$ $u=T$ $u=F$ $u=T$ $u=F$ $u=T$ $u=F$
$v=T$ 1.0 0.0 1.0 $p_{t}$ $p_{t}$ 0.0 0.0 $p_{t}$
$v=F$ 0.0 1.0 0.0 $p_{f}$ $p_{f}$ 1.0 1.0 $p_{f}$
@@ 3.5 Inferring the Posterior Probability of $Z_{0}$
Inferring the posterior probability $p(Z_{0}=T\mid\{S_{u}\})$ is a standard inference problem in hidden Markov tree models and we can apply the standard upward-downward algorithm Crouse et al. (1998) to efficiently compute this probability. For brevity, we only describe the gist of the algorithm here. More details can be found in Appendix A.4 and previous work Crouse et al. (1998); Durand et al. (2004).
The algorithm introduces an auxiliary conditional distribution $\beta(z,u)\equiv p(S_{\mathcal{T}(u)}\mid Z_{u}=z)$, where $z\in\{T,F\}$ and $S_{\mathcal{T}(u)}$ represents the confidence scores of all the nodes in the sub-tree whose root node is $u$. Note that the posterior probability of our interest, $p(Z_{0}=T\mid\{S_{\mathcal{T}(0)}\})$, can be computed as
$p(Z_{0}=T\mid\{S_{\mathcal{T}(0)}\})=\frac{\beta(T,0)p(Z_{0}=T)}{\sum_{z\in\{T,F\}}\beta(z,0)p(Z_{0}=z)},$
where $p(Z_{0}=T)$ and $p(Z_{0}=F)$ are exactly the prior probabilities we mentioned in Section 3.3, and both are set to be 0.5. Therefore, the posterior probability above can be represented as $\beta(T,0)/(\beta(F,0)+\beta(T,0))$. According to the Bayesian rule, we can derive a recursive relationship between $\beta(z,u)$ of a parent node $u$ and those of the child nodes:
$\beta(z,u)=p(S_{u}\mid Z_{u}=z)\cdot$ (1)
$\sum_{Z_{\mathcal{C}(u)}\in\{T,F\}^{m}}p(Z_{\mathcal{C}(u)}\mid Z_{u}=z)\prod_{v\in\mathcal{C}(u)}\beta(Z_{v},v),$
where $m$ is the size of $\mathcal{C}(u)$ (namely the number of child statements to node $u$). Note that the first term on the RHS is the emission probability, and the first term inside the summation is the transition probability, which are both already known. Therefore, Equation 1 provides a way to compute $\beta(z,0)$ recursively from the leaf nodes back to the root node. First, we compute the $\beta(z,u)$ of the leaf nodes as $\beta(z,u)=p(S_{u}\mid Z_{u}=z)$. Then, we use Equation 1 to recursively compute the $\beta(z,u)$ of a parent node $u$ from their child nodes, until the root node is reached. The entire process is essentially propagating and merging the beliefs in the confidence scores from sub-trees upward to the parent node. By the time we reach the root node, we have gathered the information of all the confidence scores. Hence this process is also referred to as a belief propagation process.
@@ 4 Experiments
In this section, we conduct empirical evaluations on widely-used hallucination detection benchmarks.
@@ 4.1 Experiment Configurations
We follow the previous work of hallucination detection Manakul et al. (2023); Chen et al. (2023b) and use the following datasets for evaluation: Wikibio-GPT3 Manakul et al. (2023), FELM-Science Chen et al. (2023b), and FactCheckGPT Wang et al. (2023). Wikibio-GPT3 mainly consists of biography articles generated by LLMs, whereas the other two datasets cover a broader range of topics such as physics, chemistry, and computer science.
@@
Table 3: Hallucination detection performance of different methods. We report AUROC, ROC-PR, F1 score, and detection accuracy(Acc) for all methods with two backbone models. The best results are highlighted in bold.
Evaluation settings and metrics
We conduct hallucination detection at sentence-level on Wikibio-GPT3 and FactCheckGPT and segment-level on FELM-Science following the default settings. Our evaluation metrics include the area under the receiver operator characteristic curve (AUROC), area under the precision-recall curve (AUC-PR), F1 score, and detection accuracy. As F1 score and detection accuracy evaluation require a decision threshold, we search for the optimal threshold to maximize the F1 score for each method and compute the two metrics. The hallucinated examples are considered as positive instances following the default configurations of these datasets. Moreover, Wikibio-GPT3 and FELM-Science exhibit significant class imbalances, so the detection accuracy on these datasets serve merely as reference points.
Baselines
We include the following baselines for comparison. (1) Prior Confidence , which directly queries the model’s confidence on the truthfulness of each sentence or segment. (2) CoT, which prompts the model to first generate a reasoning process before deciding the truthfulness. We adopt the prompting method from the official FELM Chen et al. (2023b) dataset, with slight modifications for sentence-level and segment-level hallucination detection. (3) SelfCheckGPT Manakul et al. (2023), which samples additional responses from the model and use the inconsistency between each response and the target statement for hallucination detection. Among the multiple variants, we choose SelfCheckGPT-prompt with the best performance for comparison. (4) Maieutic Prompting Jung et al. (2022), which builds a belief tree via backward-chaining and then infers the truth-value of the original statement that resolves the inconsistencies. (5) Semantic Uncertainty Kuhn et al. (2023); Farquhar et al. (2024), which estimate the model’s predictive uncertainty on the generation for hallucination detection. (6) Focus (Zhang et al., 2023b), which quantify the predictive uncertainty of the LLM on given texts to detect hallucinations. Note that we only include Focus when the backbone is Llama3 since it requires full access to the LLM.
Implementation details
We evaluate our methods and baselines using GPT-3.5-turbo-0125 and Llama-3-8B-Instruct. For our method, we set the maximum belief tree depth to $2$. We employ greedy decoding during belief tree construction and prior belief estimation. The exception is the statement correction strategy, where we sample $5$ corrected statements using temperature $0.7$. Additionally, since each statement in FactCheckGPT is manually processed to ensure it contains only one property or fact, we do not apply statement decomposition when build the belief tree for statements from FactCheckGPT. We use the first 120 examples in the Wikibio-GPT3 dataset to estimate the emission probability in our method, and validate it on the remaining examples and two other datasets. More details are in Appendix A.1.
4.2 Experiment Results
@@
We evaluate the effectiveness of BTProp with the experiment results in Table 3. We highlight the following observations. First, our method achieves the best performance on FELM-Science and FactCheckGPT datasets across different backbones, demonstrating the superiority of our method. BTProp improves upon the best baselines by 3% - 9% on AUROC and ROC-PR. The only exception is the Wikibio-GPT3 dataset, where the SelfCheckGPT is more effective for detection hallucinated outputs in biographies generated by LLMs. Second, compared to SelfCheckGPT which leverages contradictions between the target statement and the sampled responses for hallucination detection, our method is more effective on detecting hallucinated responses related to scientific knowledge. Both FELM-Science and FactCheckGPT datasets contain a significant proportion of questions on scientific knowledge, and our method achieves the best performance on them. Third, chain-of-thought prompting is less effective in hallucination detection, especially on the FELM-Science dataset. This finding aligns with the experimental results in the original FELM dataset paper. Our experiments show that the model tends to regard the input sentence as true most of the time, leading to sub-optimal performance. The Maieutic-prompting method, which is originall designed to verify the correctness of statements related to commonsense reasoning, is less effective for hallucination detection. Even on the FELM-Science and FactCheckGPT datasets which contain many statements about scientific knowledge, it remains less effective compared to other methods. We also find that the Semantic Uncertainty method is more effective on detection hallucinations in biography generation. However, its effectiveness diminishes when applied to the other two datasets. In contrast, our method consistently delivers competitive performance across a variety of benchmarks.
@@
One concern on our method is the belief tree construction would be inefficient. We compare the time cost of our method with the baselines on the FELM-Science and FactCheckGPT datasets, where our method significantly outperform baselines. We use the gpt-3.5-turbo as the backbone, and eliminate CoT prompting method on FELM-Science dataset due to its poor performance. We visualize the trade-off between hallucination detection performance and the time cost in Figure 4. The baseline methods, including SelfCheckGPT and CoT prompting, samples additional 20 responses for hallucination detection. We vary the number of samples and test the corresponding performance to visualize the trade-off. As depicted in the graph, the baseline performance shows diminishing returns in performance improvement with increased computational complexity. As time cost increases, the performance growth of the two baseline models gradually levels off, eventually reaching a point where additional time does not translate into significantly better performance. In comparison, our method can effectively trades off increased time cost for significantly improved performance. Despite its higher computational complexity, it achieves a marked improvement in performance metrics, underscoring its efficiency and effectiveness.
Hallucination detection in texts generated by the backbone model.
Since the datasets in the above experiments consist of text generated by another model for hallucination detection, we also include texts generated by the backbone model itself to to provide a more comprehensive evaluation of our method’s effectiveness. Detailed results are available in Appendix A.2, where we show that our method can also effectively detect the hallucinations in the text generated by the backbone model itself.
@@
We visualize several examples of the constructed belief trees to demonstrate how the inconsistencies in model’s beliefs help detect the hallucination. Due to space limit, we put these examples and the analyses in Appendix A.3.